# RoboMemory: A Brain-inspired Multi-memory Agentic Framework for Interactive Environmental Learning in Physical Embodied Systems

### v4

@@ 2.3 Closed-Loop Planning Module for Dynamic Environment
To enable closed-loop control in embodied environments, the Closed-Loop Planning Module adopts the Planner-Critic mechanism (lei2025clea, ), which consists of the planner and the critic module. For each planning step, the planner generates a long-term plan consisting of multiple steps. However, due to the dynamics of embodied environments, the action sequence in the long-term plan may become outdated during the execution of the plan. Thus, before executing each step, we use the Critic model to evaluate whether the proposed action in this step remains appropriate under the latest environment. If not, the planner will re-plan based on the latest information. The demonstration of this process is shown in Figure 2.
However, our experiments reveal that the original Planner-Critic mechanism may suffer from infinite loops. In the original mechanism, the first step of the action sequence output by the Planner is evaluated by the Critic before execution, which can lead to an infinite loop: if the Critic always demands replanning, no action will ever be executed. To address this, we modified the Planner-Critic mechanism so that the first step is not evaluated by the Critic. This ensures that even if the Critic persistently demands replanning, the RoboMemory will still execute actions.
2.4 Low-level Executor
The RoboMemory framework is a two-layer hierarchical agent framework. This design enables RoboMemory to accomplish longer-term tasks in the real world. The upper layer is responsible only for high-level planning, while the Low-level Executor carries out the actions planned by the upper layer in the real environment.
We employ a LoRA-finetuned VLA model, $\pi_{0}$ (hu2022lora, ; black2024pi_0, ), to generate manipulation actions, and a SLAM-based navigation model for locomotion. The low-level executor then translates high-level actions planned by RoboMemory into concrete arm and chassis movements in the real world.
Table 1: Comparison of Success Rates (SR) and Goal Condition Success Rates (GC) across difficulty levels between RoboMemory and baseline methods on EB-ALFRED.
Method Type Avg. Base Long
SR GC SR GC SR GC
@@ 2.4 Low-level Executor
GPT-4o Closed-source 59.0 % 68.3 % 64.0 % 74.0 % 54.0 % 62.5 %
GPT-4o-mini 17.0 % 32.4 % 34.0 % 47.8 % 0.0 % 17.0 %
Claude-3.5-Sonnet 62.0 % 63.3 % 72.0 % 72.0 % 52.0 % 54.5 %
Gemini-1.5-Pro 64.0 % 69.7 % 70.0 % 74.3 % 58.0 % 65.0 %
Gemini-2.0-flash 60.0 % 63.9 % 62.0 % 65.7 % 58.0 % 62.0 %
Llama-3.2-90B-Vision-Ins Open-source 27.0 % 33.9 % 38.0 % 43.7 % 16.0 % 24.0 %
InternVL2.5-78B 40.0 % 45.7 % 38.0 % 42.3 % 42.0 % 49.0 %
InternVL3-78B 37.0% - 38.0% - 36.0% -
InternVL2.5-38B 31.0% 36.9% 36.0% 37.3% 26.0% 36.5%
Qwen2.5-VL-72B-Ins 42.0 % - 50.0 % - 34.0 % -
@@
Voyager (Qwen2.5-VL-72B-Ins) Baselines 44.0% 63.7% 56.0% 73.2% 32.0% 54.2%
Reflexion (Qwen2.5-VL-72B-Ins) 29.0% 43.2% 48.0% 54.0% 10.0% 33.0%
Cradle (Qwen2.5-VL-72B-Ins) 43.0% 54.6% 54.0% 67.9% 32.0% 41.0%
RoboMemory (Qwen2.5-VL-72B-Ins) Ours 67.0 % 78.4 % 68.0 % 75.5 % 66.0 % 81.3 %
3 Experiments
3.1 Benchmarks
@@
To evaluate the task planning ability of RoboMemory, we select a subset of the EB-ALFRED benchmark from EmbodiedBench(yang2025embodiedbench, ). We selected the Base and Long subsets because they aim to test the agent’s planning ability. The Base and Long subset comprises 100 tasks for complex embodied tasks. The EB-ALFRED environment provides a visually-grounded operational setting that closely mimics real-world conditions (see Appendix D for environment details), enabling direct comparison with established baselines.
Additionally, we have also evaluated RoboMemory’s capabilities on the EB-Habitat benchmark (yang2025embodiedbench, ). For detailed results, please refer to Appendix B.
Moreover, we set up an environment to test the interactive environmental learning ability of RoboMemory in the real world.
3.2 Settings & Baselines
@@
To facilitate comparisons, we consider two types of baselines. First, we choose the advanced closed-source and open-source VLMs as a single agent. We compare their performance with RoboMemory. For closed source VLMs, we choose GPT-4o and GPT-4o-mini (GPT-4o-mini, ; hurst2024gpt, ), Claude3.5-Sonnet (Claude-3.5-Sonnet, ), Gemini-1.5-Pro and Gemini-2.0-flash (team2024gemini, ; Gemini2.0, ). For open source VLMs, we choose Llama-3.2-90B-Vision-Ins (llama3.2, ), InternVL-2.5-78B/28B (chen2024expanding, ), InternVL-3-72B (zhu2025internvl3, ), and Qwen2.5-VL-72B-Ins (bai2025qwen2, ). Secondly, we choose three agent frameworks: (1) Reflexion (shinn2024reflexion, ), which introduces a simple long-term memory and a self-reflection module. Reflexion uses the self-reflection module to summarize experiences as long-term memory, thereby enhancing the model’s capabilities. (2) Voyager (wang2023voyager, ), which utilizes a skill library as its procedural memory, is a widely used baseline for embodied agent planning. (3) Cradle (tan2024cradle, ), which proposes a general agent framework with episodic and procedural memory and gains good performances at various multi-model agent tasks.
@@
In our experiments, each agent framework is tested using Qwen2.5-VL-72b-Ins (qwen2.5, ). The Qwen2.5-VL-72b-Ins represents a high-performing open-source alternative. Notably, the Qwen2.5-VL-72b-Ins demonstrates performance comparable to advanced closed-source VLMs in several benchmark tasks (white2024livebench, ). We use the Qwen3-Embedding model (zhang2025qwen3, ) to create embedding vectors for RAGs in RoboMemory. For the Low-level Executor, since EB-ALFRED provides high-level action APIs, we use the low-level executor provided by EmbodiedBench instead of the VLA-based method.
@@
Where $\mathcal{X}$ denotes the test subset, and $x$ represents a test task. The success condition number ($SCN_{x}$) refers to the number of conditions the agent has accomplished, while the global condition number ($GCN_{x}$) indicates the total number of conditions required for task completion. The task is considered successful if $SCN_{x}=GCN_{x}$.
Figure 3: Efficiency improvement of Comprehensive Embodied Memory System
Table 2: Ablation Study on RoboMemory’s Success Rate (SR)
Method Avg. Base Long
@@
3.3 Main results
As shown in Table 1, our model achieves significant improvements over both single VLM agents and Agent frameworks on the EB-ALFRED. Compared to the SOTA Single VLM-Agent model, Gemini1.5-Pro, RoboMemory with Qwen2.5-VL-72B-Ins backbone improves the average SR by 3% and GC by 8.7%. This demonstrates RoboMemory’s superiority over single VLM-Agents, proving that an Agent framework with open-source models can outperform closed-source SOTA models. Furthermore, when tested against other VLM-Agent frameworks, RoboMemory also shows substantial gains. This is because, unlike other frameworks, RoboMemory’s brain-like memory system provides embodied models with more accurate and persistent contextual information. Additionally, the Planner-Critic mechanism provides a closed-loop planning ability, which helps the RoboMemory gain better performance in long-term tasks. Because the RoboMemory can detect and try to overcome possible failures. And it is more robust when encountering unexpected situations.
3.4 Efficiency Analysis
To evaluate the efficiency of the Comprehensive Embodied Memory module, we tasked RoboMemory with executing 10 long-horizon tasks, each comprising approximately 50 steps. We exclusively measured the wall-clock time consumed by memory update and retrieval operations. We analyzed the scaling behavior of memory update latency across three distinct configurations: (1) fully parallel update and retrieval across all memory modules; (2) sequential update of each memory module without parallelism; and (3) update of only the most fundamental memory component: the Temporal Memory. Results are presented in Figure 3. As shown, our parallel update strategy enables updating a multi-module memory system with latency comparable to that of updating a single base memory module. This demonstrates the critical efficiency gains afforded by parallelization across the memory architecture.
3.5 Ablation Studies
We used the full Base and Long Subset from EB-ALFRED to validate RoboMemory’s effectiveness. We removed each component systematically and observed performance changes across task categories. We use the success rate as our metric. Results are shown in Table 2.
Long-term Memory Adding long-term memory significantly improved RoboMemory’s success rate. The experiment shows that it enables interactive environmental learning while attempting to complete tasks. The semantic memory learns low-level skills’ properties, such as in what circumstances an action may fail. The temporal memory records all task attempts (successful/failed), providing valuable experience at the task level and giving insight into how to complete a task successfully. This helps the RoboMemory predict action outcomes and avoid ineffective attempts. This ability indicates that the RoboMemory has an interactive environmental learning capability.
Spatial Memory Spatial memory is crucial for embodied agents, especially given that current pretrained VLMs have limited spatial understanding ability. Our novel dynamic KG update algorithm enables KG-based spatial memory in dynamic environments. This spatial reasoning helps RoboMemory handle partially observable embodied settings.
Critic Module Table 2 shows performance without the critic module (55% vs 67% with full system). This drop highlights how the critic’s closed-loop planning adapts to dynamic environments. It helps RoboMemory recover from failures faster and handle unexpected situations better.
@@ 3.5 Ablation Studies
Figure 5: The improvement of RoboMemory after learning in the real world.
3.6 Real-world Robot Deployment
@@
To evaluate RoboMemory’s interactive environmental learning capability in the real world, we designed a kitchen environment inspired by EB-ALFRED and EB-Habitat. The scene contains 5 navigable points, 8 interactive objects, and over 10 non-interactive (but potentially distracting) items. The environment is shown in Figure 5. In the real world, we use interactive environmental video recordings captured during action execution (rather than static snapshots taken after action completion) as RoboMemory’s input. This provides a more temporally coherent perception. We created three task categories (5 tasks each). Tasks are matched to EB-ALFRED’s Base subset (avg. oracle: 10–20 steps), though actual executions often exceed 20 steps due to search and error recovery. Due to search and error recovery, the robot often exceeds 20 steps per task. Additional hardware experiment details are in Appendix D.
@@
To test the interactive environmental learning ability of RoboMemory, we run each task twice without clearing long-term memory between attempts. Success rates for first and second attempts are shown in Figure 5.
@@
The second attempt showed significantly higher success rates. This proves RoboMemory’s long-term memory effectively guides subsequent tasks in real embodied environments. Key observations include: (1) Closed-loop error recovery: RoboMemory retries failed actions when possible, even if the low-level executor (VLA model) fails. (2) Spatial reasoning: RoboMemory remembers object locations and spatial relationships using its memory. (3) Interactive environmental learning: RoboMemory analyzes failure causes reasonably. These analyses guide future decisions. Detailed examples demonstrating these capabilities and further discussions are provided in Appendix E.
@@
Moreover, we observed a significant drop in task success rates when deploying the agent with the Low-level Executor in real-world environments. This performance degradation primarily stems from the executor’s inherent limitations: (1) The VLA model exhibits unreliable instruction-following capabilities, frequently failing during grasping actions or selecting incorrect objects; (2) Pre-trained VLM models demonstrate inadequate video understanding - while capable of recognizing static objects, they struggle to interpret dynamic visual information such as action failures or state changes. These limitations collectively contribute to the reduced performance compared to simulated environments.
4 Conclusion and Future Work

### v6

@@ III-C Closed-Loop Planning Module
To enable closed-loop control in embodied environments, the Closed-Loop Planning Module adopts the Planner-Critic mechanism [25], which consists of the planner and the critic module which is powerd by same VLM model with different prompts. We denote the planner module as $\mathcal{P}$, while the critic module as $\mathcal{C}$. For each planning step, the planner generates a long-term plan consisting of multiple steps. However, due to the dynamics of embodied environments, the action sequence in the long-term plan may become outdated during the execution of the plan. Thus, before executing each step, we use the Critic model to evaluate whether the proposed action in this step remains appropriate under the latest environment. If not, the planner will re-plan based on the latest information. The demonstration of this process is shown in Figure 2.
However, our experiments reveal that the original Planner-Critic mechanism may suffer from infinite loops. In the original mechanism, the first step of the action sequence output by the Planner is evaluated by the Critic before execution, which can lead to an infinite loop: if the Critic always demands replanning, no action will ever be executed. To address this, we modify the Planner-Critic mechanism so that the first step is not evaluated by the Critic. This ensures that even if the Critic persistently demands replanning, the RoboMemory will still execute actions. The detailed algorithm is shown in Algorithm 1 in the appendix.
III-D Low-Level Executor
The RoboMemory is a two-layer hierarchical agent framework that accomplishes longer-term tasks in the real world. The upper layer is responsible only for high-level planning, while the Low-level Executor carries out the actions planned by the upper layer in the real environment.
We employ a LoRA-finetuned VLA model, $\pi_{0}$ [21, 8], to generate manipulation actions, and a SLAM-based navigation model for locomotion. The low-level executor then translates high-level actions planned by RoboMemory into concrete arm and chassis movements in the real world.
TABLE I: Comparison of Success Rates (SR) and Goal Condition Success Rates (GC) across difficulty levels (Base/Long) on EB-ALFRED and EB-Habitat benchmarks. Values are reported in percentages (%).
Method Type EB-ALFRED EB-Habitat
Average Base Long Base Long
SR GC SR GC SR GC SR GC SR GC
@@ III-D Low-Level Executor
GPT-4o Closed-source 67.0 74.9 64.0 74.0 54.0 62.5 86.0 90.7 64.0 72.2
GPT-4o-mini 30.5 40.9 34.0 47.8 0.0 17.0 74.0 77.5 14.0 21.3
Claude-3.7-Sonnet 68.5 - 68.0 - 70.0 - 90.0 - 46.0 -
Claude-3.5-Sonnet 69.5 71.8 72.0 72.0 52.0 54.5 96.0 97.5 58.0 63.3
Gemini-1.5-Pro 68.0 73.3 70.0 74.3 58.0 65.0 92.0 92.5 52.0 61.2
Gemini-2.0-flash 57.0 61.5 62.0 65.7 58.0 62.0 82.0 82.0 26.0 36.2
Llama-3.2-90B-Vision-Ins Open-source 40.5 46.6 38.0 43.7 16.0 24.0 94.0 94.5 14.0 24.3
InternVL2.5-78B 47.0 52.9 38.0 42.3 42.0 49.0 80.0 82.0 28.0 38.2
InternVL2.5-38B 37.5 42.6 36.0 37.3 26.0 36.5 60.0 61.5 28.0 35.0
InternVL3-78B 49.5 - 38.0 - 36.0 - 84.0 - 40.0 -
Qwen2.5-VL-72B-Ins 44.0 - 50.0 - 34.0 - 74.0 - 18.0 -
@@
Voyager (Qwen2.5-VL-72B-Ins) Baselines 46.5 66.4 56.0 73.2 32.0 54.2 76.0 87.0 22.0 51.0
Reflexion (Qwen2.5-VL-72B-Ins) 38.3 51.1 48.0 54.0 10.0 33.0 80.0 84.2 15.0 33.0
Cradle (Qwen2.5-VL-72B-Ins) 44.5 57.0 54.0 67.9 32.0 41.0 62.0 67.0 30.0 52.1
RoboOS (Qwen2.5-VL-72B-Ins) 25.5 33.0 32.0 38.4 12.0 17.6 38.0 47.8 20.0 28.2
RoboOS (RoboBrain2-32B) 20.0 25.7 32.0 37.2 8.0 13.2 28.0 34.8 12.0 17.4
RoboMemory (Qwen2.5-VL-72B-Ins) Ours 70.5 79.7 68.0 75.5 66.0 81.3 86.0 88.0 62.0 74.0
IV Experiments
IV-A Benchmarks
@@
To evaluate RoboMemory’s task planning ability, we select a subset of the EmbodiedBench EB-ALFRED and EB-Habitat benchmark [47]. We select the Base and Long subsets because they aim to test the agent’s planning ability. The Base and Long subsets of the two benchmarks comprise 200 tasks for complex embodied tasks. The EB-ALFRED and EB-Habitat benchmarks provide a visually grounded operational setting that closely mimics real-world conditions (see Appendix -E for environment details), enabling direct comparison with established baselines. We set temperature = 0 to avoid randomness during the experiment, aligning with the EmbodiedBench setting. Each task is executed once. Moreover, we set up an environment to test the interactive environmental learning ability of RoboMemory in the real world.
IV-B Baselines and Metrics
@@
To facilitate comparisons, we consider two types of baselines. First, we choose the advanced closed-source and open-source VLMs as a single agent. We compare their performance with RoboMemory. For closed source VLMs, we choose GPT-4o and GPT-4o-mini [31, 23], Claude3.5-Sonnet and Claude-3.7-Sonnet [2], Gemini-1.5-Pro and Gemini-2.0-flash [41, 15]. For open source VLMs, we choose Llama-3.2-90B-Vision-Ins [30], InternVL-2.5-78B/28B [11], InternVL-3-72B [55], and Qwen2.5-VL-72B-Ins [6]. Secondly, we choose three agent frameworks: (1) Reflexion [35], which introduces a simple long-term memory and a self-reflection module. Reflexion uses the self-reflection module to summarize experiences as long-term memory, thereby enhancing the model’s capabilities. (2) Voyager [45], which utilizes a skill library as its procedural memory, is a widely used baseline for embodied agent planning. (3) Cradle [39], which proposes a general agent framework with episodic and procedural memory and gains good performances at various multi-model agent tasks. (4) RoboOS [38], which proposes an embodied agent framework that consists of a Hierarchy Scene-Graph based Spatial Memory.
In our experiments, each agent framework is tested using Qwen2.5-VL-72b-Ins [42] with temperature set as 0. For the RoboOS framework, we test it on RoboBrain2-32B [40], where the RoboBrain2-32B model is designed for the RoboOS framework.
@@
The Qwen2.5-VL-72b-Ins represents a high-performing open-source alternative. Notably, the Qwen2.5-VL-72b-Ins demonstrates performance comparable to advanced closed-source VLMs in several benchmark tasks [46]. We use the Qwen3-Embedding model [53] to create embedding vectors for RAGs in RoboMemory. For the Low-level Executor, since EB-ALFRED provides high-level action APIs, we use the Low-level executor provided by EmbodiedBench instead of the VLA-based method.
@@
In above formulas, $\mathcal{X}$ denotes the test subset, and $x$ represents a test task. The success condition number ($SCN_{x}$) refers to the number of conditions the agent has accomplished, while the global condition number ($GCN_{x}$) indicates the total number of conditions required for task completion. The task is considered successful if $SCN_{x}=GCN_{x}$.
IV-C Main Results
Figure 3: Efficiency improvement of Comprehensive Embodied Memory System.
As shown in Table I, our model achieves significant improvements over both single VLM agents and Agent frameworks on the EB-ALFRED and EB-Habitat. Compared to the SOTA Single VLM-Agent model, Claude3.5-Sonnet, RoboMemory with Qwen2.5-VL-72B-Ins backbone improves the average SR by 1% and GC by 7.9%. This demonstrates RoboMemory’s superiority over single VLM-Agents, proving that an Agent framework with open-source models can outperform closed-source SOTA models. Furthermore, when tested against other VLM-Agent frameworks, RoboMemory also shows substantial gains. This is because, unlike other agent frameworks, RoboMemory’s brain-like memory system provides embodied models with more accurate and persistent contextual information. Additionally, the Planner-Critic mechanism provides a closed-loop planning ability, which helps the RoboMemory gain better performance in long-term tasks. Because the RoboMemory can detect and try to overcome possible failures. And it is more robust when encountering unexpected situations.
IV-D Efficiency Analysis
To evaluate the efficiency of the Comprehensive Embodied Memory module, we tested RoboMemory with executing 10 long-horizon tasks, each comprising approximately 50 steps. All testing is under the same hardware and software settings. We exclusively measured the wall-clock time consumed by memory update and retrieval operations. We analyzed the scaling behavior of memory update latency across three distinct configurations: (1) fully parallel update and retrieval across all memory modules; (2) sequential update of each memory module without parallelism; and (3) update of only the most fundamental memory component: the Temporal Memory. Results are presented in Figure 3. As shown, our parallel update strategy enables updating a multi-module memory system with latency comparable to that of updating a single base memory module. This demonstrates the critical efficiency gains afforded by parallelization across the memory architecture.
TABLE II: Ablation study of RoboMemory’s Success Rate (SR) on EB-ALFERD subset.
Method (EB-ALFRED) Avg. Base Long
@@ IV-D Efficiency Analysis
- w/o long-term memory 57% 66% 48%
@@
Figure 5: Real-world experiment results. “Qwen” denotes Qwen2.5-VL-72B-Ins; “RoboBrain” denotes RoboBrain2.0-32B.
IV-E Ablation Studies
We used the full Base and Long Subset from EB-ALFRED to validate RoboMemory’s effectiveness. We removed each component systematically and observed performance changes across task categories. We use the success rate as our metric. Results are shown in Table II.
Long-term Memory: Adding long-term memory significantly improves RoboMemory’s success rate. The experiment shows that it enables interactive environmental learning while attempting to complete tasks. Semantic memory learns the properties of low-level skills, such as in what circumstances an action may fail. The episodic memory records all task attempts (successful/failed), providing valuable experience at the task level and giving insight into how to complete a task successfully. This helps the RoboMemory predict the outcomes of actions and avoid ineffective attempts. This ability indicates that the RoboMemory has an interactive environmental learning capability. Additionally, we do ablations for semantic memory and episodic memory solely. The result shows that those memory modules significantly improve the capability of the performance of RoboMemory.
Spatial Memory: Spatial memory is crucial for embodied agents, especially given that current pretrained VLMs have limited spatial understanding ability. Our novel dynamic KG update algorithm enables KG-based spatial memory in dynamic environments. This spatial reasoning helps RoboMemory handle partially observable embodied settings.
Critic Module: Table II shows performance without the critic module (55% vs 67% with full system). This drop highlights how the critic’s closed-loop planning adapts to dynamic environments. It helps RoboMemory recover from failures faster and handle unexpected situations better.
IV-F Real-World Robot Deployment
@@
To evaluate RoboMemory’s interactive environmental learning capability in the real world, we designed a kitchen environment inspired by EB-ALFRED and EB-Habitat. The scene contains 5 navigable points, 8 interactive objects, and over 10 non-interactive (but potentially distracting) items. The environment is shown in Figure 5. In the real world, we use interactive environmental video recordings captured during action execution (rather than static snapshots taken after action completion) as RoboMemory’s input. This provides a more temporally coherent perception. We created three task categories (5 tasks each). Tasks are matched to EB-ALFRED’s Base subset (avg. oracle: 10–20 steps), though actual executions often exceed 20 steps due to search and error recovery. Due to search and error recovery, the robot often exceeds 20 steps per task. Additional hardware experiment details are in Appendix -E.
@@
To test the interactive environmental learning ability of RoboMemory, we ran each task twice without clearing long-term memory between attempts. Meanwhile, we compared RoboMemory against previous SOTA on real-world experiments RoboOS as baselines. The success rates for first and second attempts and different settings of RoboOS are shown in Figure 5.
@@
The second attempt showed significantly higher success rates. This proves RoboMemory’s long-term memory effectively guides subsequent tasks in real embodied environments. Key observations include: (1) Closed-loop error recovery: RoboMemory retries failed actions when possible, even if the low-level executor (VLA model) fails. (2) Spatial reasoning: RoboMemory remembers object locations and spatial relationships using its memory. (3) Interactive environmental learning: RoboMemory analyzes failure causes reasonably. These analyses guide future decisions. Detailed examples demonstrating these capabilities and further discussions are provided in Appendix -F.
@@
Moreover, we observe a significant drop in task success rates when deploying the agent with the Low-level Executor in real-world environments. This performance degradation primarily stems from the executor’s inherent limitations: (1) The VLA model exhibits unreliable instruction-following capabilities, frequently failing during grasping actions or selecting incorrect objects; (2) Pre-trained VLM models demonstrate inadequate video understanding capability. They are struggle to interpret dynamic visual information such as action failures or state changes. These limitations collectively contribute to the reduced performance compared to simulated environments.
V Conclusion